### v2

@@ B.2 Details of curve learning experiment
The first curve learning experiment in Section 4.2 aims at learning a linear function on $\mathbb{R}\to\mathbb{R}$ defined as $x\mapsto\frac{x}{2}-1$ using Leapfrog (Figure 3), Midpoint (Figure S3) and Ralston (Figure S4) ODE-net models. We take the activation function to be $f(z,\theta):=\theta_{3}\tanh(z\theta_{1}+\theta_{2})$, where $\theta_{1},\theta_{2},\theta_{3}$ denote the first, second and third coordinate of $\theta$ respectively. We also fix the output function $g$ being the identity function, and the loss function be the mean square loss (14). We initiate all models at $\theta(t):=(-1.11,0.33,1.41)$ and plotted the third coordinate of the parameter curves learned using both vanilla and modified gradient at training epochs from 0 to $15$. The training dataset consists of $128$ evenly sampled points for $x$ ranging from $2$ to $4$, and the test dataset consists of $63$ evenly sampled points for $x$ ranging from $1$ to $5$. The numbers are chosen so that we would be able to test both the interpolation and extrapolation capabilities of the trained models. We set $L=20$ for Leapfrog and $L=10$ for Midpoint and Ralston, so all of them will have $20$ stages in total. For a fair comparison, we use PyTorch’s torch.optim.SGD with identical rescaling methods applied to both inputs and labels across all experiments. We fix the random seed, use $64$ as the batch size, set the learning rate to $0.1$, and run the training process for $15$ epochs.
@@
using Leapfrog (Figure 4), Midpoint (Figure S5) and Ralston (Figure S6) ODE-net models. We use the $(x(t),y(t))$ pairs as the inputs and the corresponding time-derivatives $(\frac{dx}{dt}(t),\frac{dy}{dt}(t))$ as the labels. In each experiments, we use a randomly initiated linear mapping to bring the inputs to $\mathbb{R}^{4}$, train our ODE-net models in $\mathbb{R}^{4}$, and use another randomly initiated linear mapping to project the outputs down to $\mathbb{R}^{d}$. The two linear mappings are fixed and not trainable. We still use $\tanh$ as the activation function, and the mean square loss as the loss function. We set $L=40$ for Leapfrog and $L=20$ for Midpoint and Ralston, so all of them will have $40$ stages in total. We initiate all ODE-net layers to be the same randomly sampled parameters. The training dataset is generated by running (30) from time 0 to $5$ using Scipy’s scipy.integrate.solve_ivp function with a step size of $0.01$, so that at each time step, the state and the corresponding time derivatives yield a data/label pair. We run (30) on $[0,5]$ again with the learned models to generate the reconstructions. All optimizations are done with PyTorch’s torch.optim.SGD optimizer, fixed random seed, batch size of 64, learning rate of 0.02, and trained for 200 epochs.
@@ Acknowledgments
Appendix S1 Proof of Proposition 2.2 and 2.3
This section is dedicated to the proof of Proposition 2.2 and Proposition 2.3. In particular, Section S1.1 and Section S1.2 deal with the cases of Leapfrog and 2-stage ERK respectively, and Section S1.3 collects proof for Proposition 2.3.
S1.1 Proof of (15)
Proof S1.1.
@@
(S(1)) $\mathcal{L}(\Theta,\mathsf{z};\mathsf{p})$ $=\frac{1}{2}|g(\mathsf{z}_{L})-y|^{2}-\mathsf{p}_{\text{init}}^{\top}(\mathsf{z}_{0}-x)-\mathsf{p}_{0}^{\top}(\mathsf{z}_{1}-\mathsf{z}_{0}-hf(\mathsf{z}_{0},\mathsf{\theta}_{0}))$
@@
(S(2)) $\nabla_{\Theta}E=\frac{\partial\mathcal{L}}{\partial\Theta}+\frac{\partial\mathcal{L}}{\partial\mathsf{z}}\frac{\partial\mathsf{z}}{\partial\Theta}\,.$
If there is any $\mathsf{p}$ such that $\frac{\partial\mathcal{L}}{\partial\mathsf{z}}=0$, then $\nabla_{\Theta}E$ can be reduced to $\frac{\partial\mathcal{L}}{\partial\Theta}$, which is further characterized, according to (S(1)):
(S(3)) $\partial_{\theta_{l}}E=\frac{\partial\mathcal{L}}{\partial\mathsf{\theta}_{l}}=\mathsf{p}_{l}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
@@
(S(4)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})-\frac{\mathsf{p}_{L-1}^{\top}}{2}\,$,
(S(4)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L-1}}$ $=h\mathsf{p}_{L-1}^{\top}\partial_{z}f(\mathsf{z}_{L-1},\mathsf{\theta}_{L-1})-\frac{\mathsf{p}_{L-2}^{\top}}{2}\,$,
(S(4)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\frac{\mathsf{p}_{l+1}^{\top}}{2}+h\mathsf{p}_{l}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})-\frac{\mathsf{p}_{l-1}^{\top}}{2}\,$,
(S(4)d) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{1}}$ $=-\mathsf{p}_{0}^{\top}+h\mathsf{p}_{1}^{\top}\partial_{z}f(\mathsf{z}_{1},\mathsf{\theta}_{1})+\frac{\mathsf{p}_{2}^{\top}}{2}\,.$
@@
(S(5)) $\mathsf{p}_{L-1}^{\top}=2(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,$,
(S(6)) $\mathsf{p}_{L-2}^{\top}=2h\mathsf{p}_{L-1}^{\top}\partial_{z}f(\mathsf{z}_{L-1},\mathsf{\theta}_{L-1})\,$,
(S(7)) $\mathsf{p}_{l-1}^{\top}=\mathsf{p}_{l+1}^{\top}+2h\mathsf{p}_{l}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$,
@@
(S(8)) $\mathsf{p}_{0}^{\top}=\frac{\mathsf{p}_{2}^{\top}}{2}+h\mathsf{p}_{1}^{\top}\partial_{z}f(\mathsf{z}_{1},\mathsf{\theta}_{1})\,.$
@@
(S(9)) $\mathcal{L}(\Theta,\mathsf{z};\mathsf{p})$ $=\frac{1}{2}|g(\mathsf{z}_{L})-y|^{2}-\mathsf{p}_{0}^{\top}(\mathsf{z}_{0}-x)-\sum_{l=0}^{L-1}\mathsf{p}_{l+\alpha}^{\top}(\mathsf{z}_{l+\alpha}-\mathsf{z}_{l}-h\alpha f(\mathsf{z}_{l},\mathsf{\theta}_{l}))$
@@
therefore (S(2)) again holds true. If we can luckily find a choice of $\mathsf{p}$ so that $\frac{\partial\mathcal{L}}{\partial\mathsf{z}}=0$, then $\nabla_{\Theta}E$ is reduced to $\frac{\partial\mathcal{L}}{\partial\Theta}$, that can be easily computed, as below:
(S(10)) $\partial_{\theta_{l}}E=\frac{\partial\mathcal{L}}{\partial\mathsf{\theta}_{l}}$ $=h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$
@@
(S(11)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})-\mathsf{p}_{L}^{\top}\,$,
(S(11)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\mathsf{p}_{l+1}^{\top}+h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})-\mathsf{p}_{l}^{\top}+\mathsf{p}_{l+\alpha}^{\top}+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$,
(S(11)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l+\alpha}}$ $=\frac{h}{2\alpha}\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l+\alpha},\mathsf{\theta}_{l+\alpha})-\mathsf{p}_{l+\alpha}^{\top}\,.$
@@
(S(12)) $\mathsf{p}_{L}^{\top}=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,$,
@@
(S(13)) $\mathsf{p}_{l+\alpha}^{\top}=\frac{h}{2\alpha}\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l+\alpha},\mathsf{\theta}_{l+\alpha})\ ,$
@@
(S(14)) $\mathsf{p}_{l}^{\top}=\mathsf{p}_{l+1}^{\top}+h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})+\mathsf{p}_{l+\alpha}^{\top}+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
Combining (S(13))-(S(14)), we recover (18). Combining (S(13)) and (S(2)), and setting $\partial_{\mathsf{z}}\mathcal{L}=0$, we obtain (17).
S1.3 Proof of Proposition 2.3
@@
(S(15)) $\mathcal{L}(\Theta,\mathsf{z},\mathsf{\xi};\mathsf{p},\mathsf{q})$ $=\frac{1}{2}|g(\mathsf{z}_{L})-y|^{2}-\sum_{l=0}^{L-1}\mathsf{p}_{l+1}^{\top}(\mathsf{z}_{l+1}-\mathsf{z}_{l}-h(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})f_{l})$
@@
Similar to the previous sections, to determine the form of the cotangent vector, we ought to set $\partial_{\mathsf{z}}\mathcal{L}=0$ and $\partial_{\mathsf{\xi}}\mathcal{L}=0$. Therefore
(S(16)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})-\mathsf{p}_{L}^{\top}\,$,
(S(16)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\mathsf{p}_{l+1}^{\top}-\mathsf{p}_{l}^{\top}+\mathsf{q}_{l}^{\top}(\mathsf{e}_{\nu}\otimes\mathsf{I}_{d})$
(S(16)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{\xi}_{l}}$ $=h\mathsf{p}_{l+1}^{\top}(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})\partial_{z}f_{l}+h\mathsf{q}_{l}^{\top}(\mathsf{A}\otimes\mathsf{I}_{d})\partial_{z}f_{l}-\mathsf{q}_{l}^{\top}\,.$
@@
(S(17)) $\mathsf{p}_{L}^{\top}=(g(\mathsf{z}_{L})-y)\nabla g(\mathsf{z}_{L})\,$,
(S(18)) $\mathsf{p}_{l}^{\top}=\mathsf{p}_{l+1}^{\top}+\mathsf{q}_{l}^{\top}(\mathsf{e}_{\nu}\otimes\mathsf{I}_{d})\,$,
@@
(S(19)) $\mathsf{q}_{l}^{\top}(\mathsf{I}_{\nu d}-h(\mathsf{A}\otimes\mathsf{I}_{d})\partial_{z}f_{l})=h\mathsf{p}_{l+1}^{\top}(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})\partial_{z}f_{l}\,.$
The statement of the proposition holds by combining (S(18)) and (S(19)).
Appendix S2 Proof of Theorem 3.1
Here we present the details of the proof of Theorem 3.1. To recall the notations, we denote the solution to the ODE (1) $z(t)$, and the solution to the adjoint ODE (12) $p(t)$. The ODE-nets using (3) provides ${\mathsf{z}}_{l}$, and the auto-differentiation cotangent vector $\mathsf{p}_{l}$ satisfies (16).
S2.1 A-priori estimate
@@
This boundedness result comes from an analog to  [9] and we cite them here for the completeness of the paper. The proof extends in a straightforward way and is omitted from here. We emphasize that the boundedness is independent of the density of the discretization.
Lemma S2.2.
@@
(S(21)) $|\hat{\mathsf{p}}_{l}-{\mathsf{p}}_{l}|=\mathcal{O}(h^{2})\,.$
@@
$=2|(g(z(1))-y(x))\nabla g(z(1))-(g({\mathsf{z}}_{L})-y(x))\nabla g({\mathsf{z}}_{L})|$
@@
where we used Lemma S2.2, the Lipschitz continuity assumption on $g$ and $\nabla g$, and the boundedness of $\mathsf{z}_{L}$.
@@
where we have used the triangle inequality, and the regularity of $f$ up to its second order derivative. The last equation comes from Lemma S2.2. For $l=1,2,\dots,L-2$ we call (16) again:
@@
(S(22)) $\tilde{\mathsf{T}}:=\begin{bmatrix}I&\frac{3}{4}I&0&-\frac{1}{4}I&0&\cdots\\ \frac{1}{2}I&\frac{1}{2}I&\frac{1}{4}I&0&0&\cdots\\ 0&\frac{1}{4}I&\frac{1}{2}I&\frac{1}{4}I&0&\cdots\\ \cdots&\cdots&\ddots&\ddots&\ddots&\cdots\\ \cdots&\cdots&0&\frac{1}{4}I&\frac{1}{2}I&\frac{1}{4}I\\ \cdots&\cdots&\cdots&0&\frac{1}{4}I&\frac{1}{2}I\\ \end{bmatrix}\$,
@@
(S(23)) $|p(lh)-\hat{\mathsf{q}}_{l}|=\mathcal{O}(h^{2})\,,\quad 0\leq l\leq L-1\,$,
@@
Remark S2.7.
We recall $\mathsf{T}$ defined in (21) can also be formulated into the same form as $\tilde{\mathsf{T}}$ defined in (S(22)), with the identity matrix II being of size $n\times n$.
Proof S2.8.
@@
$\hat{\mathsf{p}}_{L-1}^{\top}=2(g(z(1))-y(x))\nabla g(z(1))=2p^{\top}(1)\,$,
@@
An analog of proof to Lemma S2.2 would show that a Leapfrog method for (12) accumulates $\mathcal{O}(h^{3})$ per iteration, and the inequality above also shows that the updates of $\hat{\mathsf{q}}$ accumulates $\mathcal{O}(h^{3})$ per iteration, so together $\Delta^{\mathsf{q}}_{l}$ collects $\mathcal{O}(h^{3})$ when updated from $l$ and $l+1$ to $l$. Upon error collection over $L=1/h$ layers, we have $|\Delta_{l}^{\mathsf{q}}|=\mathcal{O}(h^{2})$ for all $1\leq l\leq L-2$. The $\mathcal{O}$ notation contains dependence on the boundedness of $\hat{\mathsf{p}}_{l}$ and $z$, the regularity of $\theta$ up to its second order derivative, and the regularity of $f$ up to its third order derivative.
@@
Corollary S2.9 (Control of Term I in (25)).
@@
Proof S2.10.
Using Lemma S2.2 and Lemma S2.4, we obtain
@@
(S(24)) $\text{Term I}\leq\|\mathsf{T}\|_{\infty}\cdot|(L\nabla_{\Theta}E)-(L\nabla_{\Theta}^{\mathsf{p}}E)|=\mathcal{O}(h^{2})\ ,$
@@
Corollary S2.11 (Control of Term III in (25)).

### v3

@@ B.2 Details of curve learning experiment
The first curve learning experiment in Section 4.2 aims at learning a linear function on $\mathbb{R}\to\mathbb{R}$ defined as $x\mapsto\frac{x}{2}-1$ using Leapfrog (Figure 3), Midpoint (Figure SM3) and Ralston (Figure SM4) ODE-net models. We take the activation function to be $f(z,\theta):=\theta_{3}\tanh(z\theta_{1}+\theta_{2})$, where $\theta_{1},\theta_{2},\theta_{3}$ denote the first, second and third coordinate of $\theta$ respectively. We also fix the output function $g$ being the identity function, and the loss function be the mean square loss (14). We initiate all models at $\theta(t):=(-1.11,0.33,1.41)$ and plotted the third coordinate of the parameter curves learned using both vanilla and modified gradient at training epochs from 0 to $15$. The training dataset consists of $128$ evenly sampled points for $x$ ranging from $2$ to $4$, and the test dataset consists of $63$ evenly sampled points for $x$ ranging from $1$ to $5$. The numbers are chosen so that we would be able to test both the interpolation and extrapolation capabilities of the trained models. We set $L=20$ for Leapfrog and $L=10$ for Midpoint and Ralston, so all of them will have $20$ stages in total. For a fair comparison, we use PyTorch’s torch.optim.SGD with identical rescaling methods applied to both inputs and labels across all experiments. We fix the random seed, use $64$ as the batch size, set the learning rate to $0.1$, and run the training process for $15$ epochs.
@@
using Leapfrog (Figure 4), Midpoint (Figure SM5) and Ralston (Figure SM6) ODE-net models. We use the $(x(t),y(t))$ pairs as the inputs and the corresponding time-derivatives $(\frac{dx}{dt}(t),\frac{dy}{dt}(t))$ as the labels. In each experiments, we use a randomly initiated linear mapping to bring the inputs to $\mathbb{R}^{4}$, train our ODE-net models in $\mathbb{R}^{4}$, and use another randomly initiated linear mapping to project the outputs down to $\mathbb{R}^{d}$. The two linear mappings are fixed and not trainable. We still use $\tanh$ as the activation function, and the mean square loss as the loss function. We set $L=40$ for Leapfrog and $L=20$ for Midpoint and Ralston, so all of them will have $40$ stages in total. We initiate all ODE-net layers to be the same randomly sampled parameters. The training dataset is generated by running (30) from time 0 to $5$ using Scipy’s scipy.integrate.solve_ivp function with a step size of $0.01$, so that at each time step, the state and the corresponding time derivatives yield a data/label pair. We run (30) on $[0,5]$ again with the learned models to generate the reconstructions. All optimizations are done with PyTorch’s torch.optim.SGD optimizer, fixed random seed, batch size of 64, learning rate of 0.02, and trained for 200 epochs.
@@ Acknowledgments
Appendix SM1 Proof of Proposition 2.2 and 2.3
This section is dedicated to the proof of Proposition 2.2 and Proposition 2.3. In particular, Section SM1.1 and Section SM1.2 deal with the cases of Leapfrog and 2-stage ERK respectively, and Section SM1.3 collects proof for Proposition 2.3.
SM1.1 Proof of (15)
Proof SM1.1.
@@
(SM(1)) $\mathcal{L}(\Theta,\mathsf{z};\mathsf{p})$ $=\frac{1}{2}\|g(\mathsf{z}_{L})-y\|_{\mu}^{2}-\mathsf{p}_{\text{init}}^{\top}(\mathsf{z}_{0}-x)-\mathsf{p}_{0}^{\top}(\mathsf{z}_{1}-\mathsf{z}_{0}-hf(\mathsf{z}_{0},\mathsf{\theta}_{0}))$
@@
(SM(2)) $\nabla_{\Theta}E=\frac{\partial\mathcal{L}}{\partial\Theta}+\frac{\partial\mathcal{L}}{\partial\mathsf{z}}\frac{\partial\mathsf{z}}{\partial\Theta}\,.$
If there is any $\mathsf{p}$ such that $\frac{\partial\mathcal{L}}{\partial\mathsf{z}}=0$, then $\nabla_{\Theta}E$ can be reduced to $\frac{\partial\mathcal{L}}{\partial\Theta}$, which is further characterized, according to (SM(1)):
(SM(3)) $\partial_{\theta_{l}}E=\frac{\partial\mathcal{L}}{\partial\mathsf{\theta}_{l}}=\mathsf{p}_{l}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
@@
(SM(4)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}-\frac{\mathsf{p}_{L-1}^{\top}}{2}\,$,
(SM(4)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L-1}}$ $=h\mathsf{p}_{L-1}^{\top}\partial_{z}f(\mathsf{z}_{L-1},\mathsf{\theta}_{L-1})-\frac{\mathsf{p}_{L-2}^{\top}}{2}\,$,
(SM(4)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\frac{\mathsf{p}_{l+1}^{\top}}{2}+h\mathsf{p}_{l}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})-\frac{\mathsf{p}_{l-1}^{\top}}{2}\,$,
(SM(4)d) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{1}}$ $=-\mathsf{p}_{0}^{\top}+h\mathsf{p}_{1}^{\top}\partial_{z}f(\mathsf{z}_{1},\mathsf{\theta}_{1})+\frac{\mathsf{p}_{2}^{\top}}{2}\,.$
@@
(SM(5)) $\mathsf{p}_{L-1}^{\top}=2\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,$,
(SM(6)) $\mathsf{p}_{L-2}^{\top}=2h\mathsf{p}_{L-1}^{\top}\partial_{z}f(\mathsf{z}_{L-1},\mathsf{\theta}_{L-1})\,$,
(SM(7)) $\mathsf{p}_{l-1}^{\top}=\mathsf{p}_{l+1}^{\top}+2h\mathsf{p}_{l}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$,
@@
(SM(8)) $\mathsf{p}_{0}^{\top}=\frac{\mathsf{p}_{2}^{\top}}{2}+h\mathsf{p}_{1}^{\top}\partial_{z}f(\mathsf{z}_{1},\mathsf{\theta}_{1})\,.$
@@
(SM(9)) $\mathcal{L}(\Theta,\mathsf{z};\mathsf{p})$ $=\frac{1}{2}\|g(\mathsf{z}_{L})-y\|_{\mu}^{2}-\mathsf{p}_{0}^{\top}(\mathsf{z}_{0}-x)-\sum_{l=0}^{L-1}\mathsf{p}_{l+\alpha}^{\top}(\mathsf{z}_{l+\alpha}-\mathsf{z}_{l}-h\alpha f(\mathsf{z}_{l},\mathsf{\theta}_{l}))$
@@
therefore (SM(2)) again holds true. If we can luckily find a choice of $\mathsf{p}$ so that $\frac{\partial\mathcal{L}}{\partial\mathsf{z}}=0$, then $\nabla_{\Theta}E$ is reduced to $\frac{\partial\mathcal{L}}{\partial\Theta}$, that can be easily computed, as below:
(SM(10)) $\partial_{\theta_{l}}E=\frac{\partial\mathcal{L}}{\partial\mathsf{\theta}_{l}}$ $=h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{\theta}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$
@@
(SM(11)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}-\mathsf{p}_{L}^{\top}\,$,
(SM(11)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\mathsf{p}_{l+1}^{\top}+h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})-\mathsf{p}_{l}^{\top}+\mathsf{p}_{l+\alpha}^{\top}+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,$,
(SM(11)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l+\alpha}}$ $=\frac{h}{2\alpha}\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l+\alpha},\mathsf{\theta}_{l+\alpha})-\mathsf{p}_{l+\alpha}^{\top}\,.$
@@
(SM(12)) $\mathsf{p}_{L}^{\top}=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,$,
@@
(SM(13)) $\mathsf{p}_{l+\alpha}^{\top}=\frac{h}{2\alpha}\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l+\alpha},\mathsf{\theta}_{l+\alpha})\ ,$
@@
(SM(14)) $\mathsf{p}_{l}^{\top}=\mathsf{p}_{l+1}^{\top}+h(1-\frac{1}{2\alpha})\mathsf{p}_{l+1}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})+\mathsf{p}_{l+\alpha}^{\top}+h\alpha\mathsf{p}_{l+\alpha}^{\top}\partial_{z}f(\mathsf{z}_{l},\mathsf{\theta}_{l})\,.$
Combining (SM(13))-(SM(14)), we recover (18). Combining (SM(13)) and (SM(2)), and setting $\partial_{\mathsf{z}}\mathcal{L}=0$, we obtain (17).
SM1.3 Proof of Proposition 2.3
@@
(SM(15)) $\mathcal{L}(\Theta,\mathsf{z},\mathsf{\xi};\mathsf{p},\mathsf{q})$ $=\frac{1}{2}\|g(\mathsf{z}_{L})-y\|_{\mu}^{2}-\sum_{l=0}^{L-1}\mathsf{p}_{l+1}^{\top}(\mathsf{z}_{l+1}-\mathsf{z}_{l}-h(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})f_{l})$
@@
Similar to the previous sections, to determine the form of the adjoint vector, we ought to set $\partial_{\mathsf{z}}\mathcal{L}=0$ and $\partial_{\mathsf{\xi}}\mathcal{L}=0$. Therefore
(SM(16)a) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{L}}$ $=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}-\mathsf{p}_{L}^{\top}\,$,
(SM(16)b) $\frac{\partial\mathcal{L}}{\partial\mathsf{z}_{l}}$ $=\mathsf{p}_{l+1}^{\top}-\mathsf{p}_{l}^{\top}+\mathsf{q}_{l}^{\top}(\mathsf{e}_{\nu}\otimes\mathsf{I}_{d})$
(SM(16)c) $\frac{\partial\mathcal{L}}{\partial\mathsf{\xi}_{l}}$ $=h\mathsf{p}_{l+1}^{\top}(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})\partial_{z}f_{l}+h\mathsf{q}_{l}^{\top}(\mathsf{A}\otimes\mathsf{I}_{d})\partial_{z}f_{l}-\mathsf{q}_{l}^{\top}\,.$
@@
(SM(17)) $\mathsf{p}_{L}^{\top}=\langle(g(\mathsf{z}_{L})-y),\nabla g(\mathsf{z}_{L})\rangle_{\mu}\,$,
(SM(18)) $\mathsf{p}_{l}^{\top}=\mathsf{p}_{l+1}^{\top}+\mathsf{q}_{l}^{\top}(\mathsf{e}_{\nu}\otimes\mathsf{I}_{d})\,$,
@@
(SM(19)) $\mathsf{q}_{l}^{\top}(\mathsf{I}_{\nu d}-h(\mathsf{A}\otimes\mathsf{I}_{d})\partial_{z}f_{l})=h\mathsf{p}_{l+1}^{\top}(\mathsf{b}^{\top}\otimes\mathsf{I}_{d})\partial_{z}f_{l}\,.$
The statement of the proposition holds by combining (SM(18)) and (SM(19)).
Appendix SM2 Proof of Theorem 3.1
Here we present the details of the proof of Theorem 3.1. To recall the notations, we denote the solution to the ODE (1) $z(t)$, and the solution to the adjoint ODE (12) $p(t)$. The ODE-nets using (3) provides ${\mathsf{z}}_{l}$, and the auto-differentiation adjoint vector $\mathsf{p}_{l}$ satisfies (16).
SM2.1 A-priori estimate
@@
This boundedness result comes from an analog to  [DiChLiWr:2022overparameterization, Lemmas 14 and 24] and we cite them here for the completeness of the paper. The proof extends in a straightforward way and is omitted from here. We emphasize that the boundedness is independent of the density of the discretization.
Lemma SM2.2.
@@
(SM(21)) $|\hat{\mathsf{p}}_{l}-{\mathsf{p}}_{l}|=\mathcal{O}(h^{2})\,.$
@@
$=2|\langle(g(z(1))-y(x)),\nabla g(z(1))\rangle_{\mu}-\langle(g({\mathsf{z}}_{L})-y(x)),\nabla g({\mathsf{z}}_{L})\rangle_{\mu}|$
@@
where we used Lemma SM2.2, the Lipschitz continuity assumption on $g$ and $\nabla g$, and the boundedness of $\mathsf{z}_{L}$.
@@
where we have used the triangle inequality, and the regularity of $f$ up to its second order derivative. The last equation comes from Lemma SM2.2. For $l=1,2,\dots,L-2$ we call (16) again:
@@
(SM(22)) $\tilde{\mathsf{T}}:=\begin{bmatrix}I&\frac{3}{4}I&0&-\frac{1}{4}I&0&\cdots\\ \frac{1}{2}I&\frac{1}{2}I&\frac{1}{4}I&0&0&\cdots\\ 0&\frac{1}{4}I&\frac{1}{2}I&\frac{1}{4}I&0&\cdots\\ \cdots&\cdots&\ddots&\ddots&\ddots&\cdots\\ \cdots&\cdots&0&\frac{1}{4}I&\frac{1}{2}I&\frac{1}{4}I\\ \cdots&\cdots&\cdots&0&\frac{1}{4}I&\frac{1}{2}I\\ \end{bmatrix}\$,
@@
(SM(23)) $|p(lh)-\hat{\mathsf{q}}_{l}|=\mathcal{O}(h^{2})\,,\quad 0\leq l\leq L-1\,$,
@@
Remark SM2.7.
We recall $\mathsf{T}$ defined in (21) can also be formulated into the same form as $\tilde{\mathsf{T}}$ defined in (SM(22)), with the identity matrix II being of size $n\times n$.
Proof SM2.8.
@@
$\hat{\mathsf{p}}_{L-1}^{\top}=2\langle(g(z(1))-y(x)),\nabla g(z(1))\rangle_{\mu}=2p^{\top}(1)\,$,
@@
An analog of proof to Lemma SM2.2 would show that a Leapfrog method for (12) accumulates $\mathcal{O}(h^{3})$ per iteration, and the inequality above also shows that the updates of $\hat{\mathsf{q}}$ accumulates $\mathcal{O}(h^{3})$ per iteration, so together $\Delta^{\mathsf{q}}_{l}$ collects $\mathcal{O}(h^{3})$ when updated from $l$ and $l+1$ to $l$. Upon error collection over $L=1/h$ layers, we have $|\Delta_{l}^{\mathsf{q}}|=\mathcal{O}(h^{2})$ for all $1\leq l\leq L-2$. The $\mathcal{O}$ notation contains dependence on the boundedness of $\hat{\mathsf{p}}_{l}$ and $z$, the regularity of $\theta$ up to its second order derivative, and the regularity of $f$ up to its third order derivative.
@@
Corollary SM2.9 (Control of Term I in (25)).
@@
Proof SM2.10.
Using Lemma SM2.2 and Lemma SM2.4, we obtain
@@
(SM(24)) $\text{Term I}\leq\|\mathsf{T}\|_{\infty}\cdot|(L\nabla_{\Theta}E)-(L\nabla_{\Theta}^{\mathsf{p}}E)|=\mathcal{O}(h^{2})\ ,$
@@
Corollary SM2.11 (Control of Term III in (25)).